OpenStudy (hellokitty17):

Which of the following best describes the graph of f(x) = x2 - 3x - 10? Minimum at (1.5, -12.25) with intercepts at (5, 0) and (-2, 0) Minimum at (-1.5, -12.25) with intercepts at (-5, 0) and (2, 0) Minimum at (-3.5, -2.25) with intercepts at (-5, 0) and (-2, 0) Minimum at (3.5, -2.25) with intercepts at (5, 0) and (2, 0)

OpenStudy (anonymous):

i am here

OpenStudy (hellokitty17):

hey!

OpenStudy (hybrik):

To find the intersections, solve the zeros

OpenStudy (hybrik):

To do that, factor x2-3x-10

OpenStudy (hybrik):

What would you get?

OpenStudy (hellokitty17):

The second one idk

OpenStudy (hybrik):

Factor it into (x-5)(x+2)

OpenStudy (hybrik):

then solve for x in [x-5=0, x+2=0]

OpenStudy (hellokitty17):

so that first one?

OpenStudy (hellokitty17):

idk

OpenStudy (hybrik):

x=5,-2 is the zeroes.

OpenStudy (hybrik):

The minimum is the least value on Y point on ordered pair.

OpenStudy (hybrik):

Graph the equation.

OpenStudy (hybrik):

When graphing you will see a line pass through (5,0) and (-2,0)

oregonduck (oregonduck):

i will give you two medals for the answer

OpenStudy (hellokitty17):

ok

OpenStudy (hellokitty17):

@heretohelpalways

OpenStudy (anonymous):

I am here again, I don't know how to solve this so I will tag some people

oregonduck (oregonduck):

ok

OpenStudy (hellokitty17):

yes plz thanks

OpenStudy (anonymous):

@Jacob902

OpenStudy (anonymous):

@Skielerlucas04

OpenStudy (hellokitty17):

hey

OpenStudy (hybrik):

you dont need to call them, I can do this :)

OpenStudy (hellokitty17):

Help plz

oregonduck (oregonduck):

ok then do it

OpenStudy (hybrik):

Find the axis of symmetry

OpenStudy (hybrik):

-b/2a

OpenStudy (hybrik):

-3/2

OpenStudy (hellokitty17):

Just tell me the answer

OpenStudy (hybrik):

-3/2 = -1.5

OpenStudy (hybrik):

B

OpenStudy (hybrik):

i think

OpenStudy (anonymous):

lol

OpenStudy (hybrik):

no

OpenStudy (hybrik):

not b

OpenStudy (hybrik):

A

OpenStudy (hybrik):

its 100% A

OpenStudy (hybrik):

A

OpenStudy (anonymous):

ya i got A

OpenStudy (hybrik):

Medal?

OpenStudy (anonymous):

no me lmfao

OpenStudy (anonymous):

:'(

OpenStudy (hybrik):

Dont tell me your the ppl like "no u"

OpenStudy (anonymous):

huh

OpenStudy (hybrik):

thats ROBLOX

OpenStudy (hybrik):

nothing

OpenStudy (hybrik):

medal please

OpenStudy (anonymous):

nope

OpenStudy (hybrik):

I gave you sir ...

OpenStudy (anonymous):

i took that junk away

OpenStudy (hellokitty17):

medal me!

OpenStudy (anonymous):

i did

OpenStudy (hybrik):

ok and me?

OpenStudy (hybrik):

:(

OpenStudy (anonymous):

nope :P

OpenStudy (hybrik):

Im the one who did the work...

OpenStudy (anonymous):

nu uh lmao

OpenStudy (hellokitty17):

Factor completely 12a3b + 8a2b2 - 20ab3. 4ab(3a + 2ab - 5b2) 4ab(3a2 + 2ab - 5b2) 4a2b(3a3b - 2a2b2 - 5ab3) 4a2b(3a2 + 2ab - 5)

OpenStudy (hybrik):

B

OpenStudy (hellokitty17):

THANKS

OpenStudy (hybrik):

Why do I get no medals if im doing all the work?

OpenStudy (hybrik):

Great support :(